Circular No. 8762

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006jc, 2006jd, 2006je Three apparent supernovae have been found on unfiltered CCD images:

2006jc independently by K. Itagaki (Teppo-cho, Yamagata, Japan; 0.60-m f/5.7 reflector; communicated by S. Nakano, Sumoto, Japan; tabulated data below from Itagaki; additional details on CBET 666) and by T. Puckett and R. Gorelli (0.35-m reflector at Ellijay, GA; cf. IAUC 8758); and 2006jd and 2006je by R. R. Prasad and W. Li (LOSS/KAIT; cf. IAUC 8758).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006jc Oct. 9.75 9^{h}17^{m}20.^{s}78 +41^{o}54′32′′.7 13.8 11′′ W, 7′′ S
2006jd Oct. 12.54 8 02 07.43 + 0 48 31.5 17.2 22′′ E, 1′′.3 S
2006je Oct. 15.28 1 50 53.28 +33 05 53.2 17.8 19′′.2 E, 20′′.7 N

Additional magnitudes for 2006jc in UGC 4904: 2005 Dec. 11 UT, [19.8 (Puckett); 2006 Sept. 21.793, [19.5 (Itagaki); 22, [19.0 (Itagaki); Oct. 10.33, 13.8 (Puckett and Gorelli); 10.725, 13.9 (Itagaki); 11.34, 13.7 (D. Lane, Stillwater Lake, NS, 0.28-m reflector, via Puckett); 12.191, 13.9 (G. Masi and S. Foglia, Ceccano, Italy). Puckett provides position end figures 20^{s}.85, 33′′.0 for 2006jc. Itagaki adds that nothing is visible at this location on the Digitized Sky Survey (DSS). Masi provides the following position end figures for 2006jc: 20^{s}.79, 33′′.1. SN 2006jc appears to be a peculiar type-Ib/c supernova (details on CBETs 672, 674, and 677). Additional magnitudes for 2006jd in UGC 4179: Apr. 21.17, [19.0; Oct. 15.54, 17.1. Additional magnitudes for 2006je in IC 1735: Sept. 29.36, [19.0; Oct. 17.47, 18.2. SN 2006je is a peculiar type-Ia supernova (details on CBET 678).

COMET P/2006 T1 (LEVY) Elliptical orbital elements (cf. IAUC 8757) from MPEC 2006-T80:

T = 2006 Oct. 7.449 TT  ω = 179.^{o}476
e = 0.67046  Ω = 279.791 }2000.0
q = 0.98924 AU  i = 18.324

a = 3.00186 AU n^{o} = 0.189504 P = 5.20 years Visual total-magnitude estimates: Oct. 3.81 UT, 9.8 (K. Yoshimoto,

Yamaguchi, Japan, 25-cm reflector); 4.17, 9.4 (J. J. Gonzalez, Leon, Spain, 25×100 binoculars); 5.18, 10.2 (N. Biver, Versailles, France, 26-cm reflector); 9.80, 10.5 (S. Yoshida, Gunma, Japan, 40-cm reflector; full moon); 17.14, 10.0 (M. Reszelski, Szamotuly-Galowo, Poland, 13-cm reflector).

2006 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green